OpenStudy (anonymous):

find the intervals on which f is increasing \[4-x ^{2}\]

OpenStudy (anonymous):

Solve for where the solution is positive and negative. Positive = increasing, negative = decreasing.

OpenStudy (anonymous):

and how do i do that

OpenStudy (anonymous):

At this moment, I would say plug in numbers. i.e. \[4-(3)^2\] =4-9= -5 So negative and decreasing.

OpenStudy (anonymous):

You can assumed any number larger than 3 will also be negative.

OpenStudy (anonymous):

Essentially, x^2 larger than 4 will be negative but x^2 less than 4 is positive.

OpenStudy (anonymous):

ummm no that's definitely not how it has something to do with finding the derivative but idk where to go from there

OpenStudy (anonymous):

Oh! I didn't understand where you were at in your math class. In that case:

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

So set your derivative equal to 0.

OpenStudy (anonymous):

That gives you a critical number, the term you will base your intervals around.

OpenStudy (anonymous):

Different people go about the next step in different ways, but basically you want to find where the derivative is positive or negative.

OpenStudy (anonymous):

I suggest a number line: